Help Me with this Story Problem and Please Show all of the Steps. Here is the Story Problem:
A group of campers and one group leader left a campsite in a canoe traveling at a steady 8 km/h. One hour later, the other group leader left the campsite in a moterboat with all of the supplies. The moterboat followed the canoe at a steady 20km/h. How long after the canoe left the campsite did the moterboat overtake it?
That is my question.
Thanks!!
Jamie Thompson

Answer by Paul(988)

Let the time be x
The first group travel at 8mph
One hour later 'after the first' the second gruop travled at 20mph.
8(x+1)=20x
8x+8=20x
12x=8
x=8/12
x=2/3
Hence, it took them 2/3 hous to catch up to them.
Paul.